# Aligning Visual Contrastive learning models via Preference Optimization

### v1

@@ Abstract
Contrastive learning models have demonstrated impressive abilities to capture semantic similarities by aligning representations in the embedding space. However, their performance can be limited by the quality of the training data and its inherent biases. While Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO) have been applied to generative models to align them with human preferences, their use in contrastive learning has yet to be explored. This paper introduces a novel method for training contrastive learning models using Preference Optimization (PO) to break down complex concepts. Our method systematically aligns model behavior with desired preferences, enhancing performance on the targeted task. In particular, we focus on enhancing model robustness against typographic attacks, commonly seen in contrastive models like CLIP. We further apply our method to disentangle gender understanding and mitigate gender biases, offering a more nuanced control over these sensitive attributes. Our experiments demonstrate that models trained using PO outperform standard contrastive learning techniques while retaining their ability to handle adversarial challenges and maintain accuracy on other downstream tasks. This makes our method well-suited for tasks requiring fairness, robustness, and alignment with specific preferences. We evaluate our method on several vision-language tasks, tackling challenges such as typographic attacks. Additionally, we explore the model’s ability to disentangle gender concepts and mitigate gender bias, showcasing the versatility of our approach.
@@ 1 Introduction
In recent years, Vision-Language Models (VLMs) such as CLIP (Radford et al., 2021) have revolutionized downstream vision-language tasks like classification (Conde & Turgutlu, 2021), object detection (Zhong et al., 2021), segmentation (Xu et al., 2022), and image generation (Saharia et al., 2022). These models are trained on large-scale web data, such as the 400 million text-image pairs used for training CLIP. However, despite their success, VLMs exhibit several vulnerabilities. For instance, adversarial attacks (Goodfellow et al., 2015; Moosavi-Dezfooli et al., 2016) and backdoor vulnerabilities (Bai et al., 2024; Liang et al., 2024) can lead to erroneous classifications with high confidence. CLIP has also been shown to be vulnerable to typographic attacks, where text within images causes misclassification (Goh et al., 2021). Additionally, biases such as gender and racial bias (Ruggeri & Nozza, 2023; Zhang et al., 2022b; Bolukbasi et al., 2016; Jialu Wang et al., 2021) can be amplified by these models due to biases in their training datasets. Another dispreferred behavior occurs when models focus on tasks unrelated to the intended objective, as highlighted by (Menon et al., 2022). These challenges underscore the need for alignment methods in large pre-trained models.
Generative models, including Large Language Models (LLMs), have also been widely adopted for solving complex problems across various domains. Examples in the literature include GPT-4 (Achiam et al., 2023), Mistral (Albert Q. Jiang et al., 2023), and LLaMA (Touvron et al., 2023). These models are trained on massive datasets of unlabeled text, learning general language representations. Some of these LLMs require additional Supervised Fine-Tuning (SFT) to specialize for specific tasks using labeled data. However, even with fine-tuning, LLMs can produce outputs that misalign with human values or safety expectations, such as the unreliability of LLMs in autonomous driving (Chen et al., 2024). To mitigate this, alignment techniques are applied, where models are further trained using a reward model that captures human preferences (Stiennon et al., 2022; Ouyang et al., 2022b; Bai et al., 2022). Popular alignment approaches include Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022b; Christiano et al., 2023) and Direct Preference Optimization (DPO) (Rafailov et al., 2023), both of which extend the capabilities of these models beyond what SFT alone can achieve. These alignment techniques have improved the safety and relevance of LLM outputs by making them more aligned with human preferences.
@@
In this work, we extend the preference optimization paradigm to contrastive learning models, focusing on enhancing robustness against typographic attacks and mitigating gender biases. By systematically aligning the model’s behavior with human preferences, we aim to preserve its pretrained knowledge while improving its performance in sensitive areas such as fairness and robustness. The core tasks addressed in this work—mitigating typographic attacks and disentangling gender biases—underscore the versatility of our approach and its potential for improving model alignment in real-world applications. Our contributions are as the following:
  • We extend preference optimization methods, originally developed for generative models, to non-generative models, particularly contrastive learning models. This expansion opens new opportunities for improving model alignment in non-generative tasks, preserving the pretrained knowledge of models while optimizing for desired behaviors.
  • We propose controlling model behavior by adjusting the singular values of a learnable linear transformation. This allows fine-tuning of specific concepts, like gender, by modulating critical directions in the representation space while preserving overall performance. This method provides precise control and interpretability, enabling flexible post-training adjustments.
  • Our experiments demonstrate the usefulness of applying preference optimization in contrastive learning tasks, such as training for robustness against typographic attacks and disentangling gender information from the embedding space, mitigating gender bias, and showcasing its effectiveness in improving model behavior in real-world scenarios.
@@ 2 Foundations & Related Work
where $\beta>0$ controls the degree of allowed divergence from $\pi_{\text{ref}}$. Unlike RLHF, which is often slow and unstable due to its reliance on sampling and reinforcement learning, DPO (Rafailov et al., 2024) offers a more efficient and stable alternative by directly defining a loss function and bypassing reinforcement learning, allowing for standard optimization. However, DPO can sometimes overfit to preference datasets. To address this, Identity Preference Optimization (IPO) (Azar et al., 2023b) bypasses the Bradley-Terry modelization and introduces a new loss function to decrease overfitting by controlling the gap between the likelihood ratios of the model and a reference model. Defining the policy ratio as below:
@@
Additionally, newer methods like Kahneman-Tversky Optimization (KTO) (Kawin Ethayarajh et al., 2024) only require a binary signal of whether an output is desirable or undesirable, making it more practical for many real-world applications:
@@
where the weights $w(y)$ depend on whether the outcome is desirable or undesirable, and $v_{\text{KTO}}$ adjusts the model based on the reference distribution:
$v_{\text{KTO}}(x,y;\beta)=\begin{cases}\sigma(r_{\text{KTO}}(x,y)-z_{\text{ref}})&y\sim y_{\text{desirable}}|x\\ \sigma(z_{\text{ref}}-r_{\text{KTO}}(x,y))&y\sim y_{\text{undesirable}}|x\end{cases},\quad w(y)=\begin{cases}\lambda_{U}&y\sim y_{\text{undesirable}}|x\\ \lambda_{D}&y\sim y_{\text{desirable}}|x\end{cases},$
where $z_{\text{ref}}=\mathbb{E}_{x^{\prime}\sim\mathcal{D}}[\beta D_{\mathrm{KL}}(\pi_{\theta}(y^{\prime}|x^{\prime})\|\pi_{\text{ref}}(y^{\prime}|x^{\prime})]$ and $r_{\text{KTO}}(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$. Further details of mentioned methods are provided in Appendix A. These methods have been widely adopted in training generative models, including diffusion models like Denoising Diffusion Policy Optimization (DDPO) (Black et al., 2024). This work investigates how these methodologies can be applied to contrastive learning-based models. This paper focused on two critical applications: mitigating typographic attacks and addressing biases, such as gender bias in contrastive models like CLIP.
Typographic attacks. Typographic attacks exploit a model’s tendency to prioritize text over visual content, as shown by (Goh et al., 2021), causing CLIP to misclassify images based on misleading text. For example, writing ”cat” on an image of a dog could lead CLIP to incorrectly label it as ”an image of a cat”. This represents a clear example of undesired behavior in CLIP and is an ideal test case for preference optimization. Recent works, such as (Azuma & Matsui, 2023b; Joanna Materzynska et al., 2022; Ilharco et al., 2022), address this shortcoming through various fine-tuning schemes.
Gender bias, task bias, and other inductive biases. Biases such as gender and racial bias (Ruggeri & Nozza, 2023; Zhang et al., 2022b; Bolukbasi et al., 2016; Jialu Wang et al., 2021) are prevalent in vision-language models like CLIP. These models often unintentionally encode and maintain societal biases, leading to skewed predictions that disproportionately affect marginalized groups. Another common issue is task bias, where models may focus on unintended aspects of the task, further amplifying distortions in their outputs.
@@ 3 Methodology
3.1 Problem Formulation: Contrastive Learning as an MDP
3.2 Preference and Regularization Datasets
We use two datasets to facilitate this learning:
  • Regularization Dataset $\mathcal{D}_{\text{reg}}$: This dataset consists of clean samples (images or text) without adversarial manipulation. It is used during training to ensure the model retains its accuracy on clean data, enhancing its robustness in general scenarios.
By combining the preference and regularization datasets, the model is trained to align with a specific behavior or be robust against adversarial inputs while maintaining performance on clean data and other downstream tasks.
3.3 Preference-Based Contrastive Optimization
Recalling Equation 4, we first focused on policy ratio, $h_{\pi_{\theta}}(y_{w},y_{l},x)$ which we can write as the following:
Given that the objective of the model $\pi_{\theta}$ is to simultaneously increase the probability of the preferred label $\pi_{\theta}(y_{w}|x)$ and decrease the probability of the dispreferred label $\pi_{\theta}(y_{l}|x)$, the term $h_{\pi_{\theta}}(y_{w},y_{l},x)$ is positive.
Lemma 3.1
Under the assumption that the text encoder is frozen, i.e., $\mathcal{T}_{\text{ref}}=\mathcal{T}_{\theta}=\mathcal{T}$, the policy ratio for models using the contrastive learning policy, in the methods such as DPO or IPO can be expressed as:
The proof is provided in Appendix B.1.
How does the Gradient update work in case of DPO/IPO?
Corollary 3.2
Using Lemma 3.1 we can obtain the gradient of the loss as the following. Further details is provided in Appendix B.3.
@@ 3.3 Preference-Based Contrastive Optimization
Figure 1: Overview of our proposed framework, on the left side, we calculate the preference optimization loss $\mathcal{L}_{pref}(\pi_{\theta},\pi_{\text{ref}};\mathcal{D}_{\text{pref}})$ using the preference dataset, and the output logits $\mathcal{T}(y_{w})^{T}\mathcal{I}(x),\mathcal{T}(y_{l})^{T}\mathcal{I}(x)$ from both models. on the right side the regulatory loss $\mathcal{L}_{\text{reg}}(\pi_{\theta},\pi_{\text{ref}};\mathcal{D}_{\text{reg}})$ is calculated using the clean regularization dataset.
@@ 3.3.1 Regularization
In all of our training methods, we explicitly constrain our model to be proximal to the reference model on the preference dataset $\mathcal{D}_{\text{pref}}$. However, the preference dataset $\mathcal{D}_{\text{pref}}$ is often tiny and often fails to capture the entire data manifold since, in many cases, it only contains adversarial samples and out-of-distribution samples; this leads us to introduce additional regularization terms to our training, which serve to maintain proximity between the trained model and the reference.
This motivates the introduction of a regularization term $\mathcal{L}_{\text{reg}}$, designed to maintain proximity between the trained model $\pi_{\theta}$ and the reference model $\pi_{\text{ref}}$ as follow:
@@
A shortcoming from these preference losses also suggests us to introduce additional regularization terms.
From this point on, for the sake of simplicity, we denote $\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}=x_{1}$ and $\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}=x_{2}$.
In order to better understand the effect of regularization, we will build up on top of following theorem.
@@
(Feng et al., 2024) The partial derivatives of Equation 5 with respect to $x_{1}$ and $x_{2}$ are given by:
@@
We also find a similar result in the case of IPO, We leave the proof in Appendix B.2.:
@@
The partial derivatives of Equation 6 with respect to $x_{1}$ and $x_{2}$ are given by:
$\left|\frac{\partial\mathcal{L}_{\text{IPO}}(x_{1};x_{2})}{\partial x_{1}}/\frac{\partial\mathcal{L}_{\text{IPO}}(x_{1};x_{2})}{\partial x_{2}}\right|=\frac{x_{2}}{x_{1}}<1.$
In our problem formulation in section 3.1, we define the true label as $y_{w}$ and the adversarial label (e.g., typographic label) as $y_{l}$. According to Theorem 3.3 and Proposition 3.4, the rate of increase in the probability of the preferred response $y_{w}$ is lower than the rate of decrease in the probability of the dispreferred response $y_{l}$, Therefore, as suggested by Feng et al. (2024), our model primarily focuses on reducing the likelihood of the dispreferred response. This also motivates us to introduce more regulatory terms to ensure proximity.
For training, we propose using a total loss function from the dataset $\mathcal{D}=(\mathcal{D}_{\text{pref}},\mathcal{D}_{\text{reg}})$ as follows, where $\mathcal{L}_{\text{reg}}$ is the regularization loss and $\mathcal{L}_{\text{pref}}$ is the preference optimization objective:
@@
For example, consider giving images as input to CLIP, where the model chooses the corresponding caption from a fixed caption set. The selection process follows the policy we defined, using the softmax function over the similarity scores between the image and each caption in the set, as shown in Figure 1.

### v3

@@ Abstract
Contrastive learning models have demonstrated impressive abilities to capture semantic similarities by aligning representations in the embedding space. However, their performance can be limited by the quality of the training data and its inherent biases. While Preference Optimization (PO) methods such as Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO) have been applied to align generative models with human preferences, their use in contrastive learning has yet to be explored. This paper introduces a novel method for training contrastive learning models using different PO methods to break down complex concepts. Our method systematically aligns model behavior with desired preferences, enhancing performance on the targeted task. In particular, we focus on enhancing model robustness against typographic attacks and inductive biases, commonly seen in contrastive vision-language models like CLIP. Our experiments [^1] The code is available on GitHub. demonstrate that models trained using PO outperform standard contrastive learning techniques while retaining their ability to handle adversarial challenges and maintain accuracy on other downstream tasks. This makes our method well-suited for tasks requiring fairness, robustness, and alignment with specific preferences. We evaluate our method for tackling typographic attacks on images and explore its ability to disentangle gender concepts and mitigate gender bias, showcasing the versatility of our approach.
@@ 1 Introduction
In recent years, Vision-Language Models (VLMs) such as CLIP (Radford et al., 2021) have revolutionized downstream vision-language tasks like classification (Conde & Turgutlu, 2021), object detection (Zhong et al., 2021), segmentation (Xu et al., 2022), and image generation (Saharia et al., 2022). These models are trained on large-scale web data, such as the 400 million text-image pairs used for training CLIP. However, despite their success, VLMs exhibit several vulnerabilities. For instance, adversarial attacks (Goodfellow et al., 2015; Moosavi-Dezfooli et al., 2016) and backdoor vulnerabilities (Bai et al., 2024; Liang et al., 2024a) can lead to erroneous classifications with high confidence. CLIP has also been shown to be vulnerable to typographic attacks, where text within images causes misclassification (Goh et al., 2021). Additionally, biases such as gender and racial bias (Ruggeri & Nozza, 2023; Zhang et al., 2022; Bolukbasi et al., 2016; Jialu Wang et al., 2021) can be amplified by these models due to biases in their training datasets. Another dispreferred behavior occurs when models focus on tasks unrelated to the intended objective, as highlighted by (Menon et al., 2022). These challenges underscore the need for alignment methods in large pre-trained models.
Generative models, including Large Language Models (LLMs), have also been widely adopted for solving complex problems across various domains. Examples in the literature include GPT-4 (Achiam et al., 2023), Mistral (Albert Q. Jiang et al., 2023), and LLaMA (Touvron et al., 2023). These models are trained on massive datasets of unlabeled text, learning general language representations. Some of these LLMs require additional Supervised Fine-Tuning (SFT) to specialize for specific tasks using labeled data. However, even with fine-tuning, LLMs can produce outputs that misalign with human values or safety expectations, such as the unreliability of LLMs in autonomous driving (Chen et al., 2024). To mitigate this, alignment techniques are applied, where models are further trained using a reward model that captures human preferences (Stiennon et al., 2022; Ouyang et al., 2022; Bai et al., 2022). Popular alignment approaches include Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2023) and Direct Preference Optimization (DPO) (Rafailov et al., 2024), both of which extend the capabilities of these models beyond what SFT alone can achieve. These alignment techniques have improved the safety and relevance of LLM outputs by making them more aligned with human preferences.
@@
In this work, we extend the preference optimization paradigm to contrastive learning models, exemplified on enhancing robustness against typographic attacks and mitigating gender biases in image classification and retrieval. By systematically aligning the model’s behavior with human preferences, we aim to preserve its pretrained knowledge while improving its performance in sensitive areas such as fairness and robustness. The main contributions of this paper are as follows:
  • We extend and compare three PO methods, originally developed for aligning generative models, to non-generative contrastive learning models. This way, we improve model alignment, preserving the pretrained model knowledge while optimizing for desired behaviors.
  • We propose controlling model behavior by adjusting the singular values of a learnable linear transformation. This allows fine-tuning regarding specific concepts.
  • Our experiments demonstrate the benefits of applying preference optimization in contrastive learning tasks, such as training for robustness against typographic attacks and disentangling gender information from the embedding space, mitigating gender bias.
@@ 2 Foundations & Related Work
where $\beta>0$ controls the degree of allowed divergence from $\pi_{\text{ref}}$. Unlike RLHF, which is often slow and unstable due to its reliance on sampling and reinforcement learning, DPO (Rafailov et al., 2024) offers a more efficient and stable alternative by directly defining a loss function and bypassing reinforcement learning, allowing for standard optimization. However, DPO can sometimes overfit to preference datasets. To address this, Identity Preference Optimization (IPO) (Gheshlaghi Azar et al., 2024) bypasses the Bradley-Terry modelization and introduces a new loss function to decrease overfitting by controlling the gap between the likelihood ratios of the model and a reference model. Defining the policy ratio as below:
@@
Additionally, newer methods like Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) only require a binary signal of whether an output is desired or undesired, making it more practical for many real-world applications:
@@
where the weights $w(y)$ depend on whether the outcome is desired or undesired, and $v_{\text{KTO}}$ adjusts the model based on the reference distribution:
$v_{\text{KTO}}(x,y;\beta)=\begin{cases}\sigma(r_{\text{KTO}}(x,y)-z_{\text{ref}})&y\sim y_{\text{desired}}|x\\ \sigma(z_{\text{ref}}-r_{\text{KTO}}(x,y))&y\sim y_{\text{undesired}}|x\end{cases},\quad w(y)=\begin{cases}\lambda_{U}&y\sim y_{\text{undesired}}|x\\ \lambda_{D}&y\sim y_{\text{desired}}|x\end{cases},$
where $z_{\text{ref}}=\mathbb{E}_{x^{\prime}\sim\mathcal{D}}[\beta D_{\mathrm{KL}}(\pi_{\theta}(y^{\prime}|x^{\prime})\|\pi_{\text{ref}}(y^{\prime}|x^{\prime})]$ and $r_{\text{KTO}}(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$. Further details of mentioned methods are provided in Appendix A. These methods have been widely adopted in training generative models, including diffusion models like Denoising Diffusion Policy Optimization (DDPO) (Black et al., 2024). This work investigates how these methodologies can be applied to contrastive learning-based models. This paper focuses on two critical applications: mitigating typographic attacks and addressing biases, such as gender bias in contrastive visual models like CLIP.
Typographic attacks. Typographic attacks exploit a model’s tendency to prioritize text over visual content (Goh et al., 2021), causing CLIP to misclassify images based on misleading text. For example, writing ”cat” on an image of a dog could lead CLIP to incorrectly label it as ”an image of a cat”. This represents a clear example of undesired behavior in CLIP and is an ideal test case for preference optimization. Recent works, such as (Azuma & Matsui, 2023; Joanna Materzynska et al., 2022; Ilharco et al., 2022), address this shortcoming through various fine-tuning schemes.
Inductive biases. Biases such as gender and racial bias (Ruggeri & Nozza, 2023; Zhang et al., 2022; Bolukbasi et al., 2016; Jialu Wang et al., 2021) are prevalent in vision-language models that often unintentionally encode and maintain societal biases, leading to skewed predictions disproportionately affecting marginalized groups. Another common issue is task bias, where models focus on unintended aspects of the task, further amplifying distortions in their outputs (Menon et al., 2022).
@@ 3 Methodology
Figure 1 gives an overview of our approach with a joint training objective: a preference-based contrastive optimization using a dataset of preference labels $\mathcal{D}_{\text{pref}}$ and regularization using a regularization dataset $\mathcal{D}_{\text{reg}}$. Details of our methodology are given in the following.
@@
Figure 1: Overview of our proposed approach. On the left side, we calculate the preference optimization loss $\mathcal{L}_{\text{pref}}(\pi_{\theta},\pi_{\text{ref}};\mathcal{D}_{\text{pref}})$ using the preference dataset, and the output logits $\mathcal{T}(y_{w})^{T}\mathcal{I}(x),\mathcal{T}(y_{l})^{T}\mathcal{I}(x)$ from both models. On the right side, the regulatory loss $\mathcal{L}_{\text{reg}}(\pi_{\theta},\pi_{\text{ref}};\mathcal{D}_{\text{reg}})$ is calculated using the regularization dataset. The snowflake icons denote frozen encoders.
3.1 Problem Formulation: Contrastive Learning as an MDP
3.2 Preference and Regularization Datasets
We use two datasets to facilitate this learning:
  • Regularization Dataset $\mathcal{D}_{\text{reg}}$: This dataset consists of clean samples (images or text) without adversarial manipulation. It is used during training to ensure the model retains its accuracy on clean data.
By combining these datasets, the model is trained to align with a specific behavior or to be robust against adversarial inputs while maintaining performance on clean data and other downstream tasks.
3.3 Preference-Based Contrastive Optimization
Recalling Eq. (4), we first focused on policy ratio, $h_{\pi_{\theta}}(y_{w},y_{l},x)$ which we can write as:
Given that the objective of the model $\pi_{\theta}$ is to simultaneously increase the probability of the preferred label $\pi_{\theta}(y_{w}|x)$ and decrease the probability of the dispreferred label $\pi_{\theta}(y_{l}|x)$, the term $h_{\pi_{\theta}}(y_{w},y_{l},x)$ is positive. Note that $\pi_{\text{ref}}$ in this framework is kept frozen.
Lemma 3.1
Under the assumption that the text encoder is frozen, i.e., $\mathcal{T}_{\text{ref}}=\mathcal{T}_{\theta}=\mathcal{T}$, the policy ratio for models using the contrastive learning policy, in the methods such as DPO or IPO can be expressed as: [^2] See Appendix B.1 for the proof.
How does the Gradient update work in case of DPO/IPO?
Corollary 3.2
Using Lemma 3.1, we can obtain the gradient of the loss as the following. [^3] Further details are provided in Appendix B.3.
@@ 3.3.1 Regularization
In all our training methods, we explicitly constrain the model to remain proximal to the reference model on the preference dataset $\mathcal{D}_{\text{pref}}$. However, the preference dataset $\mathcal{D}_{\text{pref}}$ is typically small and often fails to capture the entire data manifold. In many cases, it primarily consists of adversarial and out-of-distribution samples. To address this limitation, we introduce additional regularization terms into our training process, ensuring the trained model maintains proximity to the reference model. This motivates the introduction of a regularization term $\mathcal{L}_{\text{reg}}$, designed to maintain proximity between the trained model $\pi_{\theta}$ and the reference model $\pi_{\text{ref}}$ as follow:
@@
A shortcoming of these preference losses also suggests introducing additional regularization terms. From this point on, for the sake of simplicity, we denote $\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}=x_{1}$ and $\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}=x_{2}$. To better understand the effect of regularization, we will build up on top of the following theorem.
@@
(Feng et al., 2024) The partial derivatives of Eq. (5) with respect to $x_{1}$ and $x_{2}$ are given by:
@@
We find a similar result in the case of IPO: [^4] See Appendix B.2 for the proof.
@@
The partial derivatives of Eq. (6) with respect to $x_{1}$ and $x_{2}$ are given by:
$\left|\frac{\partial\mathcal{L}_{\text{IPO}}(x_{1};x_{2})}{\partial x_{1}}/\frac{\partial\mathcal{L}_{\text{IPO}}(x_{1};x_{2})}{\partial x_{2}}\right|=\frac{x_{2}}{x_{1}}<1.$ (13)
In our problem formulation in Section 3.1, we define the true label as $y_{w}$ and the adversarial label (e.g., typographic label) as $y_{l}$. According to Theorem 3.3 and Proposition 3.4, the rate of increase in the probability of the preferred response $y_{w}$ is lower than the rate of decrease in the probability of the dispreferred response $y_{l}$, Therefore, as suggested by (Feng et al., 2024), our model primarily focuses on reducing the likelihood of the dispreferred response. This also motivates us to introduce more regulatory terms to ensure proximity.
Finally, for training, we propose using a total loss function from the dataset $\mathcal{D}=(\mathcal{D}_{\text{pref}},\mathcal{D}_{\text{reg}})$ as follows, where $\mathcal{L}_{\text{reg}}$ is the regularization loss and $\mathcal{L}_{\text{pref}}$ is the preference optimization objective, computed using one of Eqs. 5, 6, or 7:
@@
For example, consider giving images as input to CLIP, where the model chooses the corresponding caption from a fixed caption set. The selection process follows the policy we defined, using the softmax function over the similarity scores between the image and each caption in the set, as shown in Figure 1. The training process using the proposed loss (Eq. (14)) is outlined in Algorithm 1.
Algorithm 1 Preference-based contrastive optimization
1:dataset $\mathcal{D}=(\mathcal{D}_{\text{pref}},\mathcal{D}_{\text{reg}})$, Model $\pi_{\theta}$, Reference model $\pi_{\text{ref}}$, Regularization coef. $\lambda_{\text{reg}}$
2:for each batch $b\in\mathcal{D}$ do
3:     $b_{\text{pref}},b_{\text{reg}}\leftarrow b$triangleright Get batch of preference / regularization data
4:     Compute $\pi_{\Psi}(y|x)\triangleq\text{Softmax}(f_{\Psi}(y,x))$ for $\Psi\in\{\theta,\text{ref}\}$triangleright Compute model and ref. distributions
5:     $l_{\text{pref}}\leftarrow\mathcal{L}_{\text{po}}(\pi_{\theta},\pi_{\text{ref}};b_{\text{pref}})$triangleright Compute preference loss using one of Eqs. (5), (6), or (7)
6:     $l_{\text{reg}}\leftarrow\mathcal{L}_{\text{reg}}(\pi_{\theta},\pi_{\text{ref}};b_{\text{reg}})$triangleright Compute regularization loss as in Eq. (11), zero if $b_{\text{reg}}=\{\}$
7:     $l_{\text{tot}}\leftarrow l_{\text{pref}}+\lambda_{\text{reg}}\cdot l_{\text{reg}}$triangleright Total loss
8:     Update model $\pi_{\theta}$ by minimizing $l_{\text{tot}}$
9:end for